OpenStudy (anonymous):

Write the equation of a circle with endpoints of the diameter at (4, -3) and (-2, 5). Show your work for credit. Write the equation of a circle with endpoints of the diameter at (4, -3) and (-2, 5). Show your work for credit. @Mathematics

hero (hero):

The answer to the other one is AG = 25

OpenStudy (anonymous):

yea i kno my sister sent it to me

hero (hero):

Twiddla is a good tool to use because the files you're uploading don't work. All you have to do now is make sure to use .PNG files GIF files are the worst to use. PNG is the best

OpenStudy (anonymous):

ok will do thank u

OpenStudy (amistre64):

and your point and divide it by 2 to get your center

OpenStudy (amistre64):

(4, -3) +(-2, 5) --------- (2 , 2) then its just a matter of filling in (x-Cx)^2+(y-Cy)^2 = r^2 oh, and r = subtract the points and pythag it

OpenStudy (amistre64):

forgot to divide by 2 lol (2,2)/2 = (1,1)

OpenStudy (anonymous):

oh haha kool thank u

OpenStudy (amistre64):

(4, -3) -(-2, 5) --------- 6 , -8 ; r = sqrt(6^2 + 8^2)

OpenStudy (amistre64):

hunh, so r^2 = 6^2 + 8^2 cool

hero (hero):

You didn't find r

OpenStudy (amistre64):

i found "d" dint i

OpenStudy (amistre64):

d/2 = r ;)

hero (hero):

you removed d from dint

OpenStudy (amistre64):

(x-1)^2+(y-1)^2 = something .... should narrow it down in multiple choices lol

hero (hero):

You assume that multiple choices are available

OpenStudy (amistre64):

MCs are always available, i even include them when they are missing ;)

hero (hero):

lol

hero (hero):

Me too

OpenStudy (amistre64):

d = sqrt(6^2 + 8^2) r = d/2 = sqrt(6^2 + 8^2) r^2 = d = (6^2 + 8^2)/4

hero (hero):

Are you sure you really have to do through all that trouble just to find r?

OpenStudy (amistre64):

pretty sure, we had a meeting with the rest of my voices and it was democratically selected .... so its prolly best

OpenStudy (amistre64):

maybe do rotation of axis and such was brought up, but quickly dismissed as a bad idea

hero (hero):

You found the midpoint already

OpenStudy (amistre64):

i did?

hero (hero):

At least I thought you did

OpenStudy (amistre64):

that was the midpoint departments job. we do radius and diameter extractions over here.

hero (hero):

The the distance between the midpoint and the point on the circle would have given you the radius. You know that

OpenStudy (eyust707):

=P

OpenStudy (amistre64):

we'll make a note of it and bring it up at the next council of the head departments

OpenStudy (eyust707):

cc my department as well

OpenStudy (anonymous):

O.o

hero (hero):

I guess you have established your own council of world mathematicians

OpenStudy (amistre64):

one of thems a janitor of guatemala tho; cheaper labor costs that way

hero (hero):

Yeah, nobody wants to pay anyone big money nowadays

OpenStudy (anonymous):

why ?

OpenStudy (amistre64):

cause china is hording it all ....

OpenStudy (amistre64):

gonna have to go back to seashells and gophers teeth

OpenStudy (anonymous):

oh my

OpenStudy (amistre64):

i really should go before i wake up and find out what ive been up to ... ;)

OpenStudy (eyust707):

haha